OpenStudy (anonymous):

Fan and Medal! Rewrite the expression by using grouping symbols. y • 4 + y • z A. 4(y + z) B. y • (4 + y) • z C. z(y • 4) D. y (4 + z)

OpenStudy (acxbox22):

redo the distributive property hint 2(3+5)=2 x 3 +2 x 5

OpenStudy (anonymous):

distribution a*b+a*c=a(b+c)

OpenStudy (anonymous):

D.?

OpenStudy (acxbox22):

correct :)

OpenStudy (anonymous):

yes